# PF3plat: Pose-Free Feed-Forward 3D Gaussian Splatting

### v1

@@ Abstract
We consider the problem of novel view synthesis from unposed images in a single feed-forward. Our framework capitalizes on fast speed, scalability, and high-quality 3D reconstruction and view synthesis capabilities of 3DGS, where we further extend it to offer a practical solution that relaxes common assumptions such as dense image views, accurate camera poses, and substantial image overlaps. We achieve this through identifying and addressing unique challenges arising from the use of pixel-aligned 3DGS: misaligned 3D Gaussians across different views induce noisy or sparse gradients that destabilize training and hinder convergence, especially when above assumptions are not met. To mitigate this, we employ pre-trained monocular depth estimation and visual correspondence models to achieve coarse alignments of 3D Gaussians. We then introduce lightweight, learnable modules to refine depth and pose estimates from the coarse alignments, improving the quality of 3D reconstruction and novel view synthesis. Furthermore, the refined estimates are leveraged to estimate geometry confidence scores, which assess the reliability of 3D Gaussian centers and condition the prediction of Gaussian parameters accordingly. Extensive evaluations on large-scale real-world datasets demonstrate that PF3plat sets a new state-of-the-art across all benchmarks, supported by comprehensive ablation studies validating our design choices. Our code will be made available at: https://cvlab-kaist.github.io/PF3plat/
@@ 1 Introduction
In recent years, 3D reconstruction and novel view synthesis have garnered significant attention, particularly with the emergence of Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) and 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023). These advancements have enabled the high-quality 3D reconstruction and novel view synthesis. However, many existing methods rely on stringent assumptions, such as dense image views (Yu et al., 2024; Barron et al., 2021; 2022), accurate camera poses (Kim et al., 2022; Kwak et al., 2023; Zhu et al., 2023; Charatan et al., 2023; Chen et al., 2024), and substantial image overlaps (Yu et al., 2021; Johari et al., 2022), which limit their practical applicability. In real-world scenarios, casually captured images contain sparse and distant viewpoints, and lack precise camera poses, making it impractical to assume densely captured views with accurate camera poses. Ideally, a practical novel view synthesis solution should operate quickly and effectively with as few as two images, even under significant viewpoint changes.
@@
More recently, pose-free generalized view synthesis frameworks (Chen & Lee, 2023; Fan et al., 2023; Jiang et al., 2023; Smith et al., 2023; Hong et al., 2024) have been introduced to decouple 3D reconstruction and novel view synthesis from camera poses. Given a set of unposed images, these frameworks aim to jointly learn radiance fields and 3D geometry without relying on additional data, such as ground-truth camera pose. The learned radiance fields and geometry can then be inferred through trained neural networks, enabling single feed-forward inference. While these pioneering efforts enhance practicality, their performance remains unsatisfactory and their slow rendering speeds (Chen & Lee, 2023; Fan et al., 2023; Jiang et al., 2023; Smith et al., 2023) remain unresolved. In an attempt to boost performance and robustness under extreme scenarios when wide-baseline images are given, Hong et al. (2024) developed a unified framework that jointly estimates camera poses, correspondences, and radiance fields using additional data, such as ground truth poses, for supervision. However, this approach still inherits the inherent limitations of NeRF, including intensive memory consumption and slow rendering speeds, making it impractical for real-world applications.
In this work, we propose PF3plat (Pose-Free Feed-Forward 3D Gaussian Splatting), a novel framework for fast and photorealistic novel view synthesis from unposed images in a single feed-forward pass. Our approach leverages the efficiency and high-quality reconstruction capabilities of pixel-aligned 3DGS (Charatan et al., 2023; Szymanowicz et al., 2024), while relaxing common assumptions such as dense image views, accurate camera poses, scene-specific optimization and substantial image overlaps. However, a primary challenge in using pixel-aligned 3DGS is its dependency on precise depth and camera pose estimates for accurate localization of 3D Gaussian centers. Inaccuracies in these estimates can cause misalignments, leading to noisy or sparse gradients that destabilize training and hinder convergence, especially when the above assumptions are relaxed or scene-specific optimizations to rectify errors cannot be applied during multi-scene training.
To mitigate these issues, we find that leveraging pre-trained monocular depth estimation (Piccinelli et al., 2024) and visual correspondence (Lindenberger et al., 2023) models to achieve a coarse alignment of 3D Gaussians is highly effective, thereby promoting a stable learning process. Subsequently, we introduce learnable modules designed to refine the depth and pose estimates from the coarse alignment to enhance the quality of 3D reconstruction and view synthesis. These modules are geometry-aware and lightweight, since we leverage features from the depth network and avoid direct fine-tuning. These refined depth and pose estimates are then used to implement geometry-aware confidence scores to assess the reliability of 3D Gaussian centers, conditioning the prediction of Gaussian parameters such as opacity, covariance, and color.
@@
  • We address the challenging task of pose-free, feed-forward novel view synthesis using 3DGS, relaxing many common assumptions of existing methods to improve practicality. Our approach works well with wide-baseline images and requires no additional data beyond RGB image collections for both training and inference, highlighting its practicality.
  • To address the unique challenges arising from the misalignment of 3D Gaussians that destabilizes the learning process, we devise an approach to provide coarse alignment. We then introduce lightweight refinement modules and geometry-aware scoring functions, which not only enhance the reconstruction and view synthesis quality, but also prevent catastrophic forgetting issues typically associated with direct fine-tuning.
  • Our framework presents an effective approach to enable fast and high-performance 3D reconstruction and novel view synthesis from sparse and unposed images. We have shown that our method sets a new state-of-the-art across all benchmarks.
@@ 2 Related Work
Generalizable Scene Reconstruction and View Synthesis from Unposed Imagery.
Several innovative efforts have addressed the joint learning of camera pose and radiance fields within NeRF-based frameworks. Starting with BARF (Lin et al., 2021), subsequent research (Jeong et al., 2021; Wang et al., 2021; Bian et al., 2023; Truong et al., 2023b) has expanded upon this foundation. Notably, the use of 3D Gaussians as dynamic scene representations has led to significant advancements: Fu et al. (2023) progressively enlarges 3D Gaussians by learning transformations between consecutive frames, SplaTAM(Keetha et al., 2024) utilizes RGB-D sequences and silhouette masks to jointly update Gaussian parameters and camera poses, and InstantSplat (Fan et al., 2024) optimizes 3D Gaussians rapidly for scene reconstruction and view synthesis. Among these, methods like DBARF (Chen & Lee, 2023), FlowCAM (Smith et al., 2023), CoPoNeRF (Hong et al., 2023), and GGRt (Li et al., 2024) aim to determine camera pose and radiance fields in a single feed-forward pass. Concurrently, Splatt3R (Smart et al., 2024) builds on pre-trained 3D reconstruction models (Leroy et al., 2024) to relax some assumptions; however, it still relies on additional data such as ground truth depth and pose. In contrast, our method reconstructs 3D scenes and synthesizes novel views from unposed images with minimal overlaps in a single feed-forward pass, eliminating the need for scene-specific optimization and extensive pose data.
Monocular Depth Estimation and Correspondence Estimation.
Monocular depth estimation and visual correspondence estimation are fundamental computer vision tasks with extensive research spanning decades. Recent advancements (Yin et al., 2023; Piccinelli et al., 2024; Ke et al., 2024; Yang et al., 2024) in monocular depth estimation have matured these models, significantly enhancing various 3D vision applications. Similarly, visual correspondence estimation has advanced significantly since the emergence of deep neural networks. Conventionally, the task is evaluated primarily in camera pose estimation, where classical methods follow a pipeline that includes keypoint detection and descriptor extraction (Lowe, 2004; Bay et al., 2006; Tola et al., 2008; Salti et al., 2014; Rusu et al., 2009), tentative matching and outlier filtering (Fischler & Bolles, 1981), and camera pose estimation using solvers (Gower, 1975; Nistér, 2004; Hartley, 1997; Hesch & Roumeliotis, 2011). More recently, deep learning-based approaches have optimized each stage, outperforming traditional methods in tasks such as 2D descriptor extraction (Yi et al., 2016; DeTone et al., 2018), 3D descriptor extraction (Yew & Lee, 2018; Choy et al., 2019), sparse and dense matching (Hong & Kim, 2021; Cho et al., 2021; 2022; Hong et al., 2022a; b; 2023; Sun et al., 2021; Edstedt et al., 2024), and outlier filtering (Barath et al., 2019; Wei et al., 2023). In this work, we propose a lightweight plug-and-play depth and pose refinement modules to enhance the quality of view synthesis.
@@ 3 Method
Figure 1: Overall architecture and loss of the proposed method. (a) Given a set of unposed images, our method aligns the 3D Gaussians using a coarse-to-fine strategy. (b) In addition to photometric loss, we enforce 3D Gaussian consistency by ensuring they are placed on the same object surface through 2D-3D and 3D-3D consistency losses.
@@ 3.1 Problem Formulation
Our objective is to reconstruct a 3D scene from a set of $N$ unposed images $\{I_{i}\}_{i=1}^{N}$ with $I_{i}\in\mathbb{R}^{H\times W\times 3}$ and synthesize photo-realistic images $\hat{I}_{t}$ from novel viewpoints in a single feed-forward pass. To achieve this, we output the depth maps $\mathcal{D}_{i}\in\mathbb{R}^{H\times W}$ for each image $I_{i}$, along with their corresponding camera poses $\mathcal{P}_{i}\in\mathbb{R}^{3\times 4}$, consisting of a rotation matrix $R_{i}\in\mathbb{R}^{3\times 3}$ and a translation vector $t_{i}\in\mathbb{R}^{3\times 1}$. Additionally, we compute a set of pixel-aligned 3D Gaussians denoted as $\mathcal{G}=\{\mu_{i},\sigma_{i},\Sigma_{i},c_{i}\}_{i=1}^{N}$. Here, $\mu_{i}(p)\in\mathbb{R}^{3}$ indicates the 3D Gaussian center derived from the depth $\mathcal{D}_{i}(p)$, camera pose $\mathcal{P}_{i}$, and camera intrinsic $K_{i}$, where $p\in\mathbb{R}^{H\times W}$ represents each pixel. The opacity is represented by $\sigma_{i}(p)\in[0,1)$, $\Sigma_{i}(p)\in\mathbb{R}^{3\times 3}$ is the covariance matrix, and the color is encoded using spherical harmonics $c_{i}(p)\in\mathbb{R}^{3(L+1)}$, where $L$ is the order of the spherical harmonics. Note that, for both training and inference, we do not use any ground truth camera poses. Instead, we let the network predict the camera poses for all input images while also learning the varying scale factors across different scenes.
@@ 3.2.1 Coarse Alignment of 3D Gaussians
Inspired by recent advancements (Charatan et al., 2023; Chen et al., 2024) that highlight the advantages of pixel-aligned 3D Gaussians, such as speed, efficiency, and high-quality reconstruction and view synthesis, we extend these benefits to more challenging scenarios, specifically in the context of pose-free feed-forward view synthesis. To this end, we adopt pixel-aligned 3D Gaussians as our scene representation. However, this representation also poses certain challenges. Unlike previous methods for generalized novel view synthesis that utilize implicit representations (Chen & Lee, 2023; Smith et al., 2023; Hong et al., 2024) and benefit from the interpolation capabilities of neural networks, our approach is challenged by the explicit nature of this representation. Specifically, our method directly localizes 3D Gaussian centers using depth and camera pose estimates (Charatan et al., 2023), making it highly sensitive to inaccuracies in these estimates, which cannot be easily compensated.
Such misalignments can cause severe performance degradation and disrupt the learning process by producing sparse and noisy gradients. This issue is particularly exacerbated when wide-baseline images are given as input or the absence of ground-truth pose prevents alignments of 3D Gaussians. Without effectively addressing these challenges, we find the problem becomes nearly intractable. These issues can be mitigated by employing iterative scene-specific optimization steps or by assuming ground-truth camera poses to guide 3D Gaussians toward object surfaces. However, these solutions are incompatible with our goal of achieving a single feed-forward process from unposed images. Therefore, overcoming these limitations requires a novel strategy that can handle depth and pose ambiguities while maintaining efficiency in a feed-forward manner.
To mitigate the challenges mentioned above, we find it necessary to provide coarse alignment of 3D Gaussians, where quantitative results can be found in Tab. 4. To this end, we employ off-the-shelf models (Piccinelli et al., 2024; Lindenberger et al., 2023) to estimate initial depths $\mathcal{D}_{i}$ and camera poses $\mathcal{P}_{i}$ for our images $I_{i}$. Specifically, given depth maps $\mathcal{D}_{i}$ and sets of correspondences $\mathcal{M}_{ij}$ and their confidence values $\mathcal{C}_{ij}$ acquired from each pairwise combinations of images, e.g., $(I_{i},I_{j})$, where $i,j$ refer to image indices, we use a robust solver (Fischler & Bolles, 1981; Li et al., 2012) to estimate the relative poses $\mathcal{P}_{ij}$ between image pair. Integrating these components, we provide the necessary coarse alignment to promote stabilizing the learning process and serve as a strong foundation for further enhancements.
@@ 3.2.2 Multi-View Consistent Depth Estimation
$\begin{split}&\Delta\delta_{i}=\phi_{\mathrm{mlp}}(\mathcal{T}_{\mathrm{depth}}(F_{i})),\\ &\hat{\mathcal{D}}_{i}=\mathcal{D}_{i}+\Delta\delta_{i},\\ \end{split}$ (1)
where $\phi_{\mathrm{mlp}}(\cdot)$ is a linear projection, $\mathcal{T}_{\mathrm{depth}}$ is a deep Transformer architecture and $\Delta\delta$ is the pixel-wise depth offset. This extension promotes consistency across views and enhances performance without relying on explicit cross-attention. Instead, it leverages supervision signals derived from pixel-aligned 3D Gaussians that connect the information across views, and leverage them for the novel view synthesis task (Zhou et al., 2017) and our loss functions, which are detailed in Section 3.3. Additionally, we avoid fine-tuning the entire depth network, thereby reducing computational costs and mitigating the risk of catastrophic forgetting.

### v2

@@ Abstract
We tackle the problem of view synthesis from sparse, unposed images in a single feed-forward pass. Our method builds on 3DGS and relaxes common requirements such as dense views, accurate camera poses or depth, and large image overlaps. However, the main challenge arises from the parametrization of pixel-aligned 3D Gaussians, as their misalignments inevitably yield noisy or sparse gradients that destabilize training. To address this, we leverage pretrained monocular depth estimation and visual correspondence networks for coarse alignment, then refine depth and pose via lightweight learnable modules. We further estimate geometry confidence scores, driven by aggregated monocular and multi-view depth, to assess the reliability of 3D Gaussian centers and condition the prediction of Gaussian parameters accordingly. Extensive experiments on large-scale real-world datasets confirm that PF3plat achieves state-of-the-art performance across all benchmarks, with ablation studies validating our design choices.
Machine Learning, ICML
@@ 1 Introduction
In recent years, 3D reconstruction and view synthesis have garnered significant attention, particularly with the emergence of NeRF (Mildenhall et al., 2021) and 3DGS (Kerbl et al., 2023). These advancements have enabled high-quality 3D reconstruction and view synthesis. However, many existing methods rely on stringent assumptions, such as dense image views (Yu et al., 2024; Barron et al., 2021, 2022), accurate camera poses (Charatan et al., 2023; Chen et al., 2024), and substantial image overlaps (Yu et al., 2021; Johari et al., 2022), which limit their practical applicability. In real-world scenarios, casually captured images contain sparse and distant viewpoints, and lack precise camera poses, making it impractical to assume densely captured views with accurate camera poses.
@@
More recently, pose-free generalized view synthesis frameworks (Chen & Lee, 2023; Fan et al., 2023; Jiang et al., 2023; Smith et al., 2023; Hong et al., 2024) have been introduced to decouple view synthesis from camera poses. Given a set of unposed images, these frameworks aim to jointly learn radiance fields and 3D geometry without relying on additional data, such as ground-truth camera pose. The learned radiance fields and geometry can then be inferred through trained neural networks, enabling single feed-forward inference. While these pioneering efforts enhance practicality, their performance remains unsatisfactory and their slow rendering speeds (Chen & Lee, 2023; Fan et al., 2023; Jiang et al., 2023; Smith et al., 2023; Hong et al., 2024) remain unresolved.
In this work, we propose PF3plat (Pose-Free Feed-Forward 3D Gaussian Splatting), a novel framework for fast and photorealistic view synthesis from unposed images in a single feed-forward pass. Our approach leverages the efficiency and high-quality reconstruction capabilities of pixel-aligned 3DGS (Charatan et al., 2023; Szymanowicz et al., 2024), while relaxing common assumptions such as dense image views, accurate camera poses, scene-specific optimization and substantial image overlaps. However, a unique challenge emerges in the parametrization of pixel-aligned 3DGS. In scenarios lacking the aforementioned assumptions, inaccuracies in localizing 3D Gaussian centers can lead to noisy or sparse gradients, destabilizing training and hindering convergence. Moreover, unlike scene-specific optimization approaches (Fu et al., 2023; Fan et al., 2024), which can iteratively rectify errors at inference time, typical learning-based feed-forward frameworks are unable to benefit from such iterative refinements.
To mitigate these issues, we leverage pre-trained monocular depth estimation (Piccinelli et al., 2024) and visual correspondence (Lindenberger et al., 2023) models to achieve a coarse alignment of 3D Gaussians to promote a stable learning process. Subsequently, we introduce learnable modules designed to refine the depth and pose estimates from the coarse alignment to enhance the quality of 3D reconstruction and view synthesis. These modules are geometry-aware and lightweight, since we leverage features from the depth network and avoid direct fine-tuning. These refined depth and pose estimates are then used to implement geometry-aware confidence scores to assess the reliability of 3D Gaussian centers, conditioning the prediction of Gaussian parameters such as opacity, covariance, and color.
@@
  • We propose PF3plat, a feed-forward network that reconstructs 3D scenes, parameterized by 3D Gaussians, from sparse, unposed views without requiring ground-truth depth or pose at either training or inference phase.
  • We propose a two-stage pipeline, coarse and fine alignment, to address the unique challenges of pixel-aligned 3D Gaussian parameterization and further enhance the reconstruction and view synthesis quality, respectively.
  • Using large-scale real-world indoor and outdoor datsets, we show that our method outperforms existing works in terms of rendered image quality and the camera pose accuracy.
@@ 2 Related Work
Generalizable View Synthesis from Unposed Imagery.
Several innovative efforts have addressed the joint learning of camera pose and radiance fields within NeRF-based frameworks. Starting with BARF (Lin et al., 2021), subsequent research (Jeong et al., 2021; Wang et al., 2021; Bian et al., 2023; Truong et al., 2023b) has expanded upon this foundation. Notably, the use of 3D Gaussians as dynamic scene representations has led to significant advancements: Fu et al. (2023) progressively enlarges 3D Gaussians by learning transformations between consecutive frames, SplaTAM(Keetha et al., 2024) utilizes RGB-D sequences and silhouette masks to jointly update Gaussian parameters and camera poses, and InstantSplat (Fan et al., 2024) optimizes 3D Gaussians rapidly for scene reconstruction and view synthesis. Among these, methods like DBARF (Chen & Lee, 2023), FlowCAM (Smith et al., 2023), CoPoNeRF (Hong et al., 2023), and GGRt (Li et al., 2024) aim to determine camera pose and radiance fields in a single feed-forward pass. Concurrently, Splatt3R (Smart et al., 2024) and NoPoSplat (Ye et al., 2024) leverage pre-trained 3D reconstruction models (Leroy et al., 2024) to relax certain assumptions. However, these methods either target two-view scenarios or rely on additional supervision (e.g., ground-truth depth and pose) during training.
Monocular Depth and Correspondence Estimation
Monocular depth estimation and visual correspondence estimation are fundamental computer vision tasks that have been extensively researched over several decades. Recent advancements (Yin et al., 2023; Piccinelli et al., 2024; Ke et al., 2024; Yang et al., 2024) in monocular depth estimation have greatly matured these models, thereby expanding the capabilities of numerous 3D vision applications. Likewise, visual correspondence estimation has undergone remarkable progress since the emergence of deep neural networks. Modern approaches have optimized each step, ranging from 2D descriptor extraction (Yi et al., 2016; DeTone et al., 2018) and 3D descriptor extraction (Yew & Lee, 2018; Choy et al., 2019), to sparse and dense matching (Hong & Kim, 2021; Cho et al., 2021, 2022; Hong et al., 2022a, b, 2023; Sun et al., 2021; Edstedt et al., 2024) and outlier filtering (Barath et al., 2019; Wei et al., 2023), surpassing traditional methods in many scenarios.
Learning-based 3D Reconstruction.
Meanwhile, learning-based 3D reconstruction methods (Wang et al., 2023; Leroy et al., 2024; Wang et al., 2024) have garnered significant attention for their performance in sparse-view reconstruction, forming a robust foundation for tasks such as camera pose estimation and tracking. In contrast, our approach simultaneously tackles 3D reconstruction and view synthesis without requiring lengthy radiance field optimization at inference, enabling more direct and efficient inference.
@@ 3 Method
Figure 1: Overall architecture and loss of the proposed method. (a) Given a set of unposed images and their camera intrinsics, our method aligns the 3D Gaussians using a coarse-to-fine strategy. (b) In addition to photometric loss, we enforce 3D Gaussian consistency by ensuring they are placed on the same object surface through 2D-3D and 3D-3D consistency losses.
@@ 3.1 Problem Formulation
Our objective is to reconstruct a 3D scene from a set of $N$ unposed images $\{I_{i}\}_{i=1}^{N}$ with $I_{i}\in\mathbb{R}^{H\times W\times 3}$ and corresponding camera intrinsic $K_{i}$, to synthesize photo-realistic images $\hat{I}_{t}$ from novel viewpoints in a single feed-forward pass. Note that, in line with existing pose-free view synthesis methods (Fu et al., 2023; Ye et al., 2024; Hong et al., 2024; Chen & Lee, 2023; Smith et al., 2023), we assume camera intrinsic parameters are available, as they are generally available from modern devices (Arnold et al., 2022). To render, we output the depth maps $\mathcal{D}_{i}\in\mathbb{R}^{H\times W}$ for each image $I_{i}$, along with their corresponding camera poses $\mathcal{P}_{i}\in\mathbb{R}^{3\times 4}$, consisting of a rotation matrix $R_{i}\in\mathbb{R}^{3\times 3}$ and a translation vector $t_{i}\in\mathbb{R}^{3\times 1}$. Additionally, we compute a set of pixel-aligned 3D Gaussians denoted as $\mathcal{G}=\{\mu_{i},\sigma_{i},\Sigma_{i},c_{i}\}_{i=1}^{N}$. Here, $\mu_{i}(p)\in\mathbb{R}^{3}$ indicates the 3D Gaussian center derived from the depth $\mathcal{D}_{i}(p)$, camera pose $\mathcal{P}_{i}$, and camera intrinsic $K_{i}$, where $p\in\mathbb{R}^{H\times W}$ represents each pixel. Finally, the opacity is represented by $\sigma_{i}(p)\in[0,1)$, $\Sigma_{i}(p)\in\mathbb{R}^{3\times 3}$ is the covariance matrix, and the color is encoded using spherical harmonics $c_{i}(p)\in\mathbb{R}^{3(L+1)}$, where $L$ is the order of the spherical harmonics.
@@ 3.2.1 Coarse Alignment of 3D Gaussians
Despite numerous advantages such as speed, efficiency, and high-quality reconstruction and view synthesis (Charatan et al., 2023; Chen et al., 2024), pixel-aligned 3D Gaussians poses certain challenges. Unlike previous methods for generalized novel view synthesis that utilize implicit representations (Chen & Lee, 2023; Smith et al., 2023; Hong et al., 2024) and benefit from the interpolation capabilities of neural networks, our approach is challenged by the explicit nature of this representation. Specifically, inaccuracies in localizations of 3D Gaussian centers makes it highly vulnerable to noisy and sparse graidents, which cannot be easily compensated.
Such misalignments and sprase gradients can either cause severe performance degradation or disrupt the learning process. This issue is particularly exacerbated when wide-baseline images are given as input or the absence of ground-truth pose or depth prevents alignments of 3D Gaussians. Without effectively addressing these challenges, we find the problem becomes nearly intractable, as we demonstrate in Tab. 4. A possible solution to mitigate this issue is to empoloy iterative scene-specific optimization steps (Fu et al., 2023) or to assume ground-truth camera poses or depth as a guidance for stable learning process (Ye et al., 2024). However, these solutions are incompatible with our goal of achieving a single feed-forward process with training solely from unposed images. Therefore, overcoming these limitations requires a novel strategy that can handle depth and pose ambiguities while maintaining efficiency in a feed-forward manner.
To this end, we propose to provide coarse alignment of 3D Gaussians. We employ off-the-shelf models (Piccinelli et al., 2024; Lindenberger et al., 2023) to estimate initial depths $\mathcal{D}_{i}$ and camera poses $\mathcal{P}_{i}$ for our images $I_{i}$, while other variants can also be leveraged, as we show in supplementary material. Specifically, given depth maps $\mathcal{D}_{i}$ and sets of correspondences $\mathcal{M}_{ij}$ and their confidence values $\mathcal{C}_{ij}$ acquired from each pairwise combinations of images, e.g., $(I_{i},I_{j})$, where $i,j$ refer to image indices, we use a robust solver (Fischler & Bolles, 1981; Li et al., 2012) to estimate the relative poses $\mathcal{P}_{ij}$ between image pairs. Integrating these components, we provide the necessary coarse alignment to promote stabilizing the learning process and serve as a strong foundation for further enhancements.
@@ 3.2.2 Multi-View Consistent Depth Estimation
$\Delta\delta_{i}=\phi_{\mathrm{mlp}}(\mathcal{T}_{\mathrm{depth}}(F_{i})),\ \ \ \ \ \ \hat{\mathcal{D}}_{i}=\mathcal{D}_{i}+\Delta\delta_{i},\$ (1)
where $\phi_{\mathrm{mlp}}(\cdot)$ is a linear projection, $\mathcal{T}_{\mathrm{depth}}$ is a deep Transformer architecture and $\Delta\delta_{i}$ is the pixel-wise depth offset. This extension promotes consistency across views and enhances performance without relying on explicit cross-attention. Instead, it leverages supervision signals derived from pixel-aligned 3D Gaussians that connect the information across views, and leverage them for the novel view synthesis task (Zhou et al., 2017) and our loss functions, which are detailed in Section 3.3. Additionally, we avoid fine-tuning the entire depth network, thereby reducing computational costs and mitigating the risk of catastrophic forgetting.